Prev Next

SW&IT Meeting Minutes 2020-01-27

Attendees: Carl, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200113
LR20200113 - Bec, Matthieu 72.0
LR20200113 - Holmberg, Carl 40.0
LR20200113 - Colson, Andrew 72.0
LR20200113 - Edmeades, David 72.0
LR20200113 - Sola, Igor 0.0
LR20200113 - Kubanek, Petr 0.0
LR20200113 - Hooper, Stephen 72.0
LR20200113 - Zhang, Yang 69.0
LR20200113 - Zhang, Xianyu 78.0

News from management

  • S/W support schedule 2020A updated (Feb-Jul) YZ joining in, PK, AC leaving out.

Previous Action Items

Owner Action Due Date Completion Date
DE Summit Veam Backup (soul-telemetry, etc) + bring back VM2 until PEPSI/VM arrives 2020-01-20  
DE sciarc webinar details from UITS asap, CV sent invite for Feb 5  

Open topics

  • OSA Nightly Report 2020-01-25 - azcam/gcs ideas for improvements (IS)
  • Summer Shutdown proposals for planning (all)
  • Office rework, new leds - your choice of color temperature (DE)

Team reports

MB

  • Last week
    • OC, WPC, Management
    • phone interviews
    • DMS websocket/mongodb service
    • job post
  • This week
    • GMT review Monday
    • DMS continued
    • telemetry conclusion
    • resume and conclude AO data with XZ
    • hydrogen AHS
    • TCS ITs review with PK and YZ

DE

  • Veeam quote in process, will follow up today
  • Called Academic Tech this morning re: Webinar
    • Fronline is poking the ticket owner; anticipate movement today
    • Not clear on why this is hoop-jumpy
  • Racked PEPSI servers and burned in over the weekend
  • Julian's keychain gone again; may have to rebuild
  • Made Olga happy with NFS tools install
  • Synology fixes
    • Failed drives replaced
      • WD/Hitachi weirdness; had to send image of drive label to support to get an RMA since the serial was rejected by the online forms
    • Google seems to be coming down on the "less secure apps" thing; modified the Synology config to use the aspmx.l.google.com SMTP server
  • UITS seems to be unplugging Ethernet ports in low-use rooms (555, 527). Had techs out to fix.
  • Lighting requests? I am in contact with Steve Lee from Electrical Services.

XZ

  • Last week
    • Zoom call with MG and Guido about power backplane
    • PB machine setup with Stephen
    • DX ASM bump monitoring
    • AO meeting
  • This week
    • Power Backplane
    • attend SOUL commission

YZ

  • Developed a MODS file checksum script.
  • Some TCS review and planning.
  • Finished all current safety training.
  • Vacation until Feb 5.

SH

  • No more dental stuff for at least a month

week before MLK:

  • office moves from fifth to third
  • Xianyu call about VPN
  • met with Paul G. about his rsync
  • telemetry work, cleanup and discussion

last week:
  • more office moves on Tuesday morning
  • Cleaning core filters (Leroy), debug CONEX ethernet for Dan (with Leroy)
  • AGW2 help
  • Light up AO machine in lab for PB
  • DNS servers. No cutover yet, but working on it.
  • RPM for weather station stuff installed on obs1, robs, rm507-{4..5}

next:
  • particle monitor
  • UAPD webcam
  • ASC debugging
  • DNS cutover in Tucson
  • Telemetry jiggering from a couple of weeks ago.

CH

Last week
  • AGW: Assisted Alessandro C and John Morris with getting AGW2 aligned via azcamsvr.
  • AHS: experimented with draft INDI driver.
  • DMS: extended imonnit data ingest to include TMS laser power.
  • ALS: began work on laser alert cron job. Polls imonnit.com for laser power sensor reading.

This week
  • AGW: Advise John Morris on getting xfreerdp/MS Remote Desktop running on his Mac so he can directly access azcamsvr.
  • AHS: Test LBTI INDI driver draft, screen scraping vacuum reading. Query Paul G on proper API to query for this data.
  • ALS: laser power cron job to write to EPICS TMS location if laser power reading over 0.258V.
  • DMS: experiment with websockets to stream Mongo query results in soft real time.

IS

Last week:
  • Time Accounting Tool
    • Adding editing to the Queue Time Accounting tool
    • Adding new backup and poorweather program tabs and change logic to accrue to the general Backup and Poorweather bins in the Downtime tab.

This week:
  • Time Accounting Tool
    • Find a way to reroute in the server and the client the "reporting" part.
    • Export to excel
    • Look into django admin login capabilities for the tool

Topic revision: r14 - 03 Feb 2020, IgorSola
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback